# COOL-LAMPS. VII. Quantifying Strong-lens Scaling Relations with 177 Cluster-scale Gravitational Lenses in DECaLS

### v3

@@ Abstract
We compute parametric measurements of the Einstein-radius-enclosed total mass for 177 cluster-scale strong gravitational lenses identified by the ChicagO Optically-selected Lenses Located At the Margins of Public Surveys (COOL-LAMPS) collaboration with lens redshifts ranging from $0.2\lessapprox z\lessapprox 1.0$ using only two measured parameters in each lensing system: the Einstein radius, and the brightest-cluster-galaxy (BCG) redshift. We then constrain the Einstein-radius-enclosed luminosity and stellar mass by fitting parametric spectral energy distributions (SEDs) with aperture photometry from the Dark Energy Camera Legacy Survey (DECaLS) in the $g$, $r$, and $z$-band Dark Energy Camera (DECam) filters. We find that the BCG redshift, enclosed total mass, and enclosed luminosity are strongly correlated and well described by a planar relationship in 3D space. We also find that the enclosed total mass and stellar mass are correlated with a logarithmic slope of $0.443\pm 0.035$, and the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope of $-0.563\pm 0.035$. The correlations described here can be used to validate strong-lensing candidates in upcoming imaging surveys—such as Rubin/Legacy Survey of Space and Time (LSST)—in which an algorithmic treatment of lensing systems will be needed due to the sheer volume of data these surveys will produce.
Galaxy clusters (584) — High-redshift galaxy clusters (2007) — Scaling relations (2031) — Spectral energy distribution (2129) — Strong gravitational lensing (1643)
@@ 1 Introduction
Strong gravitational lensing is a rare phenomenon in the universe in which an intervening object with sufficient surface mass density bends light from a background source relative to the observer such that multiple images of the source are formed. For cases where the background source is a galaxy, the lensing effect can create highly distorted and magnified images of the source galaxy in an arclike shape (e.g., Lynds & Petrosian 1986; Soucail et al. 1987; Kneib & Natarajan 2011; Meneghetti et al. 2013; Rivera-Thorsen et al. 2017; Shajib et al. 2022). The discovery of new gravitational lenses of this type is accelerating rapidly, particularly with the help of machine learning applied to finding galaxy-galaxy strong lenses in large optical-imaging surveys (e.g., Huang et al. 2021; Rojas et al. 2022; Zaborowski et al. 2023). Coupled with access to a flood of new data from large surveys such as Rubin/LSST in the coming years, the population of candidate lenses both in the bulk and at the margins will only grow. Contemporary methodologies often rely on visual inspection in order to winnow candidate lists to an acceptable number worthy of detailed follow-up (e.g., Diehl et al. 2017); such inspection has been successful in locating new systems thus far (e.g., Rojas et al. 2023). However, with thousands to millions of lines of sight to choose from in an ever-growing plethora of imaging data, statistical measures and correlations in existing lensing systems may help hone candidate-lens samples by quantifying instances of strong lensing in tandem with traditional morphological identification. This would produce a purer sample and help streamline the ability to identify targets. In this paper, we analyze a sample of 177 strong gravitational lenses identified via visual inspection by the COOL-LAMPS collaboration in DECaLS Legacy Survey Data Release 8 (LS DR8, Dey et al. 2019) images, and we investigate the scaling relations between the BCG redshift, Einstein-radius-enclosed (enclosed) total mass, enclosed luminosity, enclosed stellar mass, and enclosed stellar-to-total mass fraction therein. The sample we use here consists of a refined subsample of all strong-lensing candidates identified by COOL-LAMPS, and we consider them unambiguous and robust as informed by their visual morphology—despite a lack of definite spectroscopic confirmation.
@@
Throughout this paper, we adopt a $\mathrm{\Lambda CDM}$ WMAP-9 cosmology (Hinshaw et al., 2013). All photometric calculations were done in the AB magnitude system.
@@ 2 Methodology
Galaxies and galaxy clusters host their mass in a variety of forms. Theory and observations both place strong constraints on observed stellar luminosity as a function of mass (e.g., Kuiper 1938; Presotto et al. 2014; Wang & Zhong 2018), hot intracluster gas contributes significantly to the total mass in galaxy clusters (e.g., Forman et al. 1972; Kellogg et al. 1972; Markevitch & Vikhlinin 2007), and both galaxies and galaxy clusters are composed of a significant fraction of dark matter (e.g., Zwicky 1933; Rubin 1986; Navarro et al. 1996; Bryan & Norman 1998; Hradecky et al. 2000). However, robustly measuring these mass components on a per-system basis is time and resource intensive—which is unsustainable for large samples. Conveniently, the existence of lensing prescribes a specific amount of mass that must exist to create the observed lensing a priori.
@@
Remolina González et al. (2020) have shown that knowing only the angle between the BCG and lensed source arc (a proxy for Einstein radius; $\theta_{E}$) and the line-of-sight geometry (redshift of the lens and the source; $z_{L}$, $z_{S}$), the core mass interior to the Einstein radius of a strong-lensing galaxy cluster may be measured with minimal scatter on order of $\approx 10\%$. With these three parameters, the mass can be immediately calculated and accounts for all of the projected mass density interior to the Einstein radius. Measuring mass in this way also offers a well-defined aperture with which to constrain photometric properties of the lens as well. By directly integrating over pixels inside this aperture in flux-calibrated images, the lens-galaxy photometry can be automatically utilized by an SED modeller to obtain measurements of enclosed luminosity and stellar mass.
@@
The COOL-LAMPS collaboration has constructed a large dataset of strong-lensing candidates using a simple human-led ranking system (e.g., Khullar et al. 2021; Sukay et al. 2022; Martinez et al. 2023; Zhang et al. 2023; Napier et al. 2023b; Klein et al. 2024). From this, we have collated a sample of 177 galaxy clusters and groups that exhibit robust visual evidence of strong gravitational lensing in which the primary lensed source arc is roughly circular with its projected center located at or very near to the BCG. Any systems with complex source-galaxy morphology are less likely to be well described by the simple symmetric dark-matter profiles that inform the parametric mass estimates described in Remolina González et al. (2020)—such complex systems were intentionally excluded from this work.
Figure 1: A subset of nine strong gravitational lenses analyzed in this work. The lenses are arranged by increasing redshift from the top to the bottom of the figure, and they are arranged by increasing Einstein radius from the left to the right of the figure. In each subplot, red circles denote positional constraints used to derive the Einstein radius, white circles represent the Einstein aperture, and a black ‘x’ marks the exact coordinates of the BCG. Respective lens redshifts and Einstein radii are also shown in each subplot. RGB images have equal astrometric scaling and are sourced from DECaLS LS DR9 (Dey et al., 2019) using $z$, $r$, and $g$-band imaging data with custom color scaling respectively.
Figure 2: Observed-frame spectral energy distributions (SEDs) in maggies for the nine lenses shown in Figure 1 after SED fitting with Prospector. The top plot of each subplot contains the best-fit SED (red line), 1024 SEDs drawn from their respective posterior distributions (grey lines), observed lens-galaxy photometry (blue points), and photometric filter transmission curves for the Dark Energy Camera (DECam, Flaugher et al. 2015) $g$, $r$, and $z$-band filters from left to right respectively (in arbitrary units). The bottom plot of each subplot shows the residual between the observed photometry and the best-fit SED—normalized by the standard deviation of the observed photometry.
@@ 2.1 Einstein Radius
We defined the Einstein radius for each system as the radius of a circle centered on the BCG (Remolina González et al., 2020) which minimizes the total angular separation between each of the three most readily identifiable bright “clumps” in the tangentially lensed arcs and the perimeter of said circle. Henceforth, we refer to this circular aperture centered on the BCG with radius equal to the Einstein radius as the “Einstein aperture”. Four points (the BCG center and three points along each arc) physically constrain the Einstein aperture, and we obtained them by tagging the point of peak surface brightness for the BCG and similar peaks in surface brightness in the clumps of each tangentially lensed arc respectively. If three distinct peaks were unable to be found along the tangential arcs, the main arc was simply traced. We bootstrapped the three tangential-arc positional constraints 1000 times, and we adopted the mean and standard deviation from the resulting distribution of Einstein radii as the Einstein radius and its error respectively for each system. A descriptive visualization sampling the range of Einstein radii and lens redshifts considered in this work is shown in Figure 1.
@@ 2.2 Lens Redshift
BCGs are sui generis red-sequence cluster galaxies which can also be used as proxies to infer properties of the cluster-scale dark-matter halo (e.g., Oegerle & Hoessel 1991; Lauer et al. 2014). Since we assume that the lensing systems in this work can be well described by a symmetric dark-matter profile, we take the redshift of the BCG as interchangeable with the systematic redshift of the cluster-scale dark-matter halo. The lens redshift converts the measured Einstein radius to a proper transverse distance via the angular diameter distance, and we queried the Sloan Digital Sky Survey Data Release 15 (SDSS DR15, Aguado et al. 2019; Kollmeier et al. 2019) in tandem with the Legacy Survey Data Release 9 (LS DR9, Dey et al. 2019; Duncan 2022) for each system in this work to obtain lens (BCG) redshifts. If the centroid coordinates for each BCG corresponded to a warning-free spectroscopic redshift in SDSS DR15, we adopted that redshift as the lens redshift for that system along with its associated error. If a given BCG lacked any corresponding value in SDSS DR15, as is the case for the higher redshift clusters in this work and/or clusters located in SDSS-non-imaged areas, we adopted the LS DR9 photometric redshift (Zhou et al., 2023) as the lens redshift and its associated error instead. We obtained spectroscopic redshifts for 41 systems and photometric redshifts for the remaining 136 systems.
@@
Y $t_{\mathrm{age}}$ Age of the cluster in Gyr. Top Hat: [$t_{z=20}$, $t_{z_{BCG}}$].
Y $\tau$ SFH e-folding time in Gyr. Top Hat: [0.1, 10].
@@ 2.3 Source Redshift
In strong lensing, the source redshift is typically constrained either with spectroscopic follow-up (e.g., Sharon et al. 2020) or by inferring a photometric redshift (e.g., Cerny et al. 2018). However, obtaining these redshifts for all identified source galaxies in a given system requires prolonged additional inquiry in tension with the appeal of an efficient mass estimator. Remolina González et al. (2020) have shown that substituting a single known source-galaxy redshift with a distribution of source-galaxy redshifts introduces a statistically insignificant uncertainty into the final distribution of mass measurements when compared to the magnitude of other systematic uncertainties. This distribution of lensed source redshifts is measurable (e.g., Bayliss et al. 2011a, b; Tran et al. 2022), and we simply adopt a well-described Gaussian with $\mu=2$ and $\sigma=0.2$ from Bayliss et al. (2011a) in keeping with the methodology of Remolina González et al. (2020) as the distribution of source redshifts across the entire sample in this work.

### v4

# COOL-LAMPS. VII. Quantifying Strong-lens Scaling Relations with 177 Cluster-scale Strong Gravitational Lenses in DECaLS

@@ Abstract
We estimate the Einstein-radius-enclosed total mass for 177 cluster-scale strong gravitational lenses identified by the ChicagO Optically selected Lenses Located At the Margins of Public Surveys (COOL-LAMPS) collaboration with lens redshifts ranging from $0.2\lessapprox z\lessapprox 1.0$ using the brightest-cluster-galaxy (BCG) redshift and an observable proxy for the Einstein radius. We constrain the Einstein-radius-enclosed luminosity and stellar mass by fitting parametric spectral energy distributions to aperture photometry from the Dark Energy Camera Legacy Survey in the $g$-, $r$-, and $z$-band Dark Energy Camera filters. We find that the BCG redshift, enclosed total mass, and enclosed luminosity are strongly correlated and well described by a planar relationship in 3D space. We find that the enclosed total mass and stellar mass are correlated with a logarithmic slope of $0.500^{+0.029}_{-0.031}$, and the enclosed total mass and stellar-to-total mass fraction are correlated with a logarithmic slope of $-0.495^{+0.032}_{-0.033}$. In tandem with the small radii within which these slopes are constrained, this may suggest invariance in baryon conversion efficiency and feedback strength as a function of cluster-centric radii in galaxy clusters. Additionally, the correlations described here should have utility in ranking strong-lensing candidates in upcoming imaging surveys—such as Rubin/Legacy Survey of Space and Time—in which an algorithmic treatment of strong lenses will be needed due to the sheer volume of data these surveys will produce.
Galaxy clusters (584) ; High-redshift galaxy clusters (2007) ; Scaling relations (2031) ; Spectral energy distribution (2129) ; Strong gravitational lensing (1643)
journal: ApJfacilities: Sloan (BOSS spectrograph), Blanco (DECam wide-field camera)software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), astro-prospector (Johnson et al., 2021), emcee (Foreman-Mackey et al., 2013), FSPS (Conroy et al., 2009; Conroy & Gunn, 2010), Jupyter Notebook (Kluyver et al., 2016), LENSTOOL (Jullo et al., 2007), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (McKinney, 2010), python-FSPS (Johnson et al., 2023), SAOImageDS9 (Joye & Mandel, 2003), scipy (Virtanen et al., 2020).
@@ 1 Introduction
Figure 1: A normalized histogram of the fitted Einstein radii in log space for this work shown in blue. Overplotted in black is a normalized Gaussian with one-sided Gaussian roll-off fit to the cumulative distribution function (CDF) of these radii in dex (see Section 2.1 for parameters describing this fit). Also plotted in red is a Gaussian probability distribution function with $\mu=0.73$ and $\sigma=0.316$ from Zitrin et al. (2012) representing the simulated Einstein radius distribution of 10,000 Sloan Digital Sky Survey (SDSS) clusters.
@@
Strong gravitational lensing is a rarely observed phenomenon in the Universe in which an intervening object with sufficient surface-mass density bends light from a background source relative to the observer such that multiple images of the source are formed. For cases where the background source is a galaxy, the lensing effect can create highly distorted and magnified images of the source galaxy in an arc-like shape (e.g., Lynds & Petrosian 1986; Soucail et al. 1987; Kneib & Natarajan 2011; Meneghetti et al. 2013; Rivera-Thorsen et al. 2017; Shajib et al. 2024). The discovery of new strong lenses is accelerating rapidly, particularly with the help of machine learning applied to large optical-imaging surveys (e.g., Huang et al. 2021; Rojas et al. 2022; Zaborowski et al. 2023). Coupled with access to a flood of new data from large surveys such as Rubin/Legacy Survey of Space and Time (LSST) in the coming years, the population of candidate lenses both in the bulk and at the margins will only grow.
Any flux-limited survey will have tentative arcs; i.e. at the flux limit. Regardless of the lensing configuration, numerous faint features that are suggestive of strong lensing, but whose reality and interpretation as strong lensing, will exist and require verification. Contemporary methodologies of identifying strong lenses often rely on visual inspection in order to winnow candidate lists to an acceptable number worthy of detailed follow-up (e.g., Diehl et al. 2017); such inspection has been successful in locating new systems thus far (e.g., Rojas et al. 2023). In addition, automated strong-lens modeling has emerged as a promising avenue toward blindly identifying multiple-image families in strong lenses (e.g., Stapelberg et al. 2019; Carrasco et al. 2020; Zalesky & Ebeling 2020), though such methodologies typically rely on a light-traces-mass model to feed to a preliminary lens model—of which the latter can be computationally expensive. With thousands to millions of lines of sight to choose from in an ever-growing plethora of imaging data, analyzing statistical measures and correlations in existing strong lenses may help hone candidate-lens samples by quantifying instances of cluster-scale strong lensing in tandem with geometric identification.
Figure 2: A subset of nine strong gravitational lenses analyzed in this work. The lenses are arranged by increasing redshift from the top to the bottom of the Figure, and they are arranged by increasing Einstein radius from the left to the right of the Figure. Respective lens redshifts and Einstein radii are also shown in each subplot. RGB images have equal astrometric scaling and are sourced from DECaLS LS DR9 (Dey et al., 2019) using $z$-, $r$-, and $g$-band imaging data with custom color scaling, respectively. This Figure is available as an animation. The animation fades in and out red circles, which denote positional constraints used to derive the Einstein radius, white circles representing the Einstein aperture, and a black ‘x’ marking the exact coordinates of each BCG. (An animation of this figure is available in the online article.)
This work aims to quantify scaling relations in existing cluster-scale strong lenses such that ambiguous systems could be analyzed with a similar methodology to determine if a given system lies on or off the scaling relationships we derive here. Coupled with existing algorithms that can identify preferential lines of sight for cluster-scale strong lensing (e.g., Wong et al. 2013) and an arc-finding algorithm (e.g., Seidel & Bartelmann 2007), these calculations could be undertaken automatically for arbitrary imaging data. This analysis would help streamline the ability to identify targets for follow-up observations by ranking candidate lenses to produce a purer sample. We analyze a sample of 177 cluster-scale strong gravitational lenses identified via visual inspection by the COOL-LAMPS collaboration in DECaLS Legacy Survey Data Release 8 (LS DR8, Dey et al. 2019) images, and we investigate the scaling relations between the brightest-cluster-galaxy (BCG) redshift, Einstein-radius-enclosed (enclosed) total mass, enclosed luminosity, enclosed stellar mass, and enclosed stellar-to-total mass fraction therein. The sample we use here consists of a refined subsample of all strong-lensing candidates identified by COOL-LAMPS that we consider robust—despite a lack of definite spectroscopic confirmation in many cases. All of the lenses used here have the visual geometry of a giant arc with clear curvature that is bright enough to be unambiguously visible in the survey data.
Figure 3: Observed-frame spectral energy distributions (SEDs) in maggies for the nine lenses shown in Figure 2 after SED fitting with Prospector. The top plot of each subplot contains the best-fit SED (red line), 1024 SEDs drawn from their respective posterior distributions (gray lines), observed cluster photometry (blue points), and photometric filter transmission curves for the Dark Energy Camera (DECam; Flaugher et al. 2015) $g$-, $r$-, and $z$-band filters from left to right, respectively (in arbitrary units). The bottom plot of each subplot shows the residual between the observed photometry and the best-fit SED—normalized by the standard deviation of the observed photometry.
@@
Throughout this paper, we adopt a flat cold dark matter WMAP-9 cosmology (Hinshaw et al., 2013). All photometric calculations were done in the AB magnitude system.
@@ 2 Methodology
Galaxies and galaxy clusters host their mass in a variety of forms. Theory and observations both place strong constraints on observed stellar luminosity as a function of mass (e.g., Kuiper 1938; Presotto et al. 2014; Wang & Zhong 2018), hot intracluster gas and diffuse intracluster light contribute to the total mass in galaxy clusters (e.g., Forman et al. 1972; Kellogg et al. 1972; Gonzalez et al. 2005; Markevitch & Vikhlinin 2007), and both galaxies and galaxy clusters are composed of a significant fraction of dark matter (e.g., Zwicky 1933; Rubin 1986; Navarro et al. 1996; Bryan & Norman 1998; Hradecky et al. 2000). However, robustly measuring these mass components on a per-system basis is time and resource intensive—which is unsustainable for large samples. Conveniently, the existence of lensing prescribes a specific amount of mass that must exist to create the observed lensing a priori.
@@
Remolina González et al. (2020) have shown that the core mass interior to the Einstein radius of a strong-lensing galaxy cluster may be computed with an uncertainty of $\approx 10\%$ via the simple evaluation of a function parameterized by the angle between the BCG and lensed source arc (a proxy for Einstein radius; $\theta_{E}$) and the line-of-sight geometry (redshift of the lens and the source; $z_{L}$, $z_{S}$). With these three parameters, the mass can be immediately calculated and accounts for all of the projected mass density interior to the Einstein radius for a strong lens whose dark matter halo can be described by a circularly symmetric dark matter profile. Measuring mass in this way also offers a well-defined aperture with which to constrain photometric properties of the strong lens as well. By directly integrating over pixels that correspond to cluster members inside this aperture in flux-calibrated images, the photometry of the strong lens can be automatically utilized by a spectral energy distribution (SED) modeler to obtain measurements of enclosed luminosity and stellar mass.
@@
The COOL-LAMPS collaboration has constructed and analyzed a large dataset of strong-lensing candidates using a simple human-led ranking system (e.g., Khullar et al. 2021; Sukay et al. 2022; Martinez et al. 2023; Zhang et al. 2023; Napier et al. 2023b; Klein et al. 2024; Cloonan et al. 2024). From this, we have collated a sample of 177 galaxy clusters that exhibit robust visual evidence of strong gravitational lensing in which the primary lensed source arc is roughly circular with its projected center located at or very near to the BCG. Any systems with complex source-galaxy geometry in the image plane imply that the lens possesses a complex dark matter distribution that is less likely to be well described a circularly symmetric dark matter profile. Since the assumption of a circularly symmetric dark matter profile underpins the functions used to estimate the enclosed mass described in Remolina González et al. (2020), such complex systems were intentionally excluded from this work.
@@ 2.1 Einstein Radius
We defined the Einstein radius for each system as the radius of a circle centered on the BCG (Remolina González et al., 2020) that minimizes the total angular separation between each of the three most readily identifiable bright “clumps” in the tangentially lensed arcs and the perimeter of said circle. Henceforth, we refer to this circular aperture centered on the BCG with radius equal to the Einstein radius as the “Einstein aperture”. Four points (the BCG center and three points along each arc) physically constrain the Einstein aperture, and we obtained them by tagging the point of peak surface brightness for the BCG and similar peaks in surface brightness in the clumps of each tangentially lensed arc, respectively. If three distinct peaks were unable to be found along the tangential arcs, the main arc was simply traced. Each tangential-arc point was assigned a Gaussian uncertainty in both spatial axes with $\mu$ equal to its location and $\sigma=0.5^{\prime\prime}$ which roughly corresponds to the width of an average arc. We bootstrapped the three tangential-arc points 1000 times, with each sampling drawn from the 2D Gaussian distribution for each point. We adopted the mean and standard deviation from the resulting distribution of Einstein radii as the Einstein radius and its error, respectively, for each system. By fitting to the cumulative distribution function (CDF) of the Einstein radii for all 177 systems in this work, we find that they are distributed log normally with $\mu=0.71\pm 0.07$ and $\sigma=0.31\pm 0.04$. We also parameterize an incompleteness roll-off to low values with a one-sided Gaussian centered at $0.49\pm 0.16$ with $\sigma=0.14\pm 0.07$ which indicates that our sample is incomplete starting at $3.09^{\prime\prime}\,{}^{+1.37}_{-0.99}$ and 50% incomplete by $2.11^{\prime\prime}\,{}^{+1.08}_{-0.74}$. Physical motivations for this incompleteness are further discussed in Section 5.2. Zitrin et al. (2012) simulated the “universal” distribution of Einstein radii for a sample of 10,000 Sloan Digital Sky Survey (SDSS) clusters and found that this distribution is log normal with $\mu=0.73^{+0.02}_{-0.03}$ and $\sigma=0.316^{+0.004}_{-0.002}$. As shown in Figure 1, we find good agreement between the distribution of Einstein radii in this work and that of Zitrin et al. (2012)—indicating that our methodology is a good proxy for the true underlying Einstein radius for our sample of galaxy clusters.
@@ 2.2 Lens Redshift
BCGs are sui generis red-sequence cluster galaxies that can also be used as proxies to infer properties of the cluster-scale dark matter halo (e.g., Oegerle & Hoessel 1991; Lauer et al. 2014). Since we assume that the strong lenses in this work can be well described by a symmetric dark matter profile, we take the redshift of the BCG as interchangeable with the systematic redshift of the cluster-scale dark matter halo. The lens redshift converts the measured Einstein radius to a proper transverse distance via the angular diameter distance, and we queried the SDSS Data Release 15 (DR15, Aguado et al. 2019; Kollmeier et al. 2019) in tandem with the Legacy Survey Data Release 9 (LS DR9, Dey et al. 2019; Duncan 2022) for each system in this work to obtain lens (BCG) redshifts. If the centroid coordinates for each BCG corresponded to a warning-free spectroscopic redshift in SDSS DR15, we adopted that redshift as the lens redshift for that system along with its associated error. If a given BCG lacked any corresponding value in SDSS DR15, as is the case for the higher-redshift clusters in this work and/or clusters located in SDSS-non-imaged areas, we adopted the LS DR9 photometric redshift (Zhou et al., 2023) as the lens redshift and its associated error instead. We obtained spectroscopic redshifts for 41 systems and photometric redshifts for the remaining 136 systems. A descriptive visualization sampling the range of Einstein radii and lens redshifts considered in this work is shown in Figure 2.
@@
Y $t_{\mathrm{age}}$ Age of the cluster in gigayears. Top Hat: [$t_{z=20}$, $t_{z_{BCG}}$].
Y $\tau$ SFH e-folding time in gigayears. Top Hat: [0.1, 10].
@@ 2.3 Source Redshift
In strong lensing, the source redshift is typically constrained either with spectroscopic follow-up (e.g., Sharon et al. 2020) or by inferring a photometric redshift (e.g., Cerny et al. 2018). However, obtaining these redshifts for all identified source galaxies in a given system requires prolonged additional inquiry in tension with the appeal of an efficient mass estimator. Remolina González et al. (2020) showed that substituting a single known source-galaxy redshift with a distribution of source-galaxy redshifts introduces a statistically insignificant uncertainty into the final distribution of mass measurements when compared to the magnitude of other systematic uncertainties. This distribution of lensed source redshifts is measurable (e.g., Bayliss et al. 2011a, b; Tran et al. 2022), and we simply adopt a well-described Gaussian with $\mu=2$ and $\sigma=0.2$ from Bayliss et al. (2011a) in keeping with the methodology of Remolina González et al. (2020) as the distribution of source redshifts across the entire sample in this work. We need not consider the redshift distribution of sources other than giant arcs, since all of the source galaxies we analyze in this work are geometrically consistent with the giant arcs analyzed in Bayliss et al. (2011a). While we note that a later paper by Bayliss (2012) favors a slightly smaller error with the same mean, using this distribution would only marginally impact the uncertainties on the calculated total masses. We therefore use the distribution in Bayliss et al. (2011a) to maintain continuity with Remolina González et al. (2020).